OpenStudy (anonymous):

Compute Triangle DEF with vertices D (-4, -5) E (-1, 3) F (4, 4)

OpenStudy (anonymous):

ok i mean the area if someone didnt get what compute it ment :P so the area of this triangle with these vetices HELP PLEASE D:

OpenStudy (anonymous):

Area = (1/2)|y1(x2 - x3) + y2(x3 - x1) + y3(x1 - x2)| = (1/2)|x1(y2 - y3) + x2(y3 - y1) + x3(y1 - y2)| so you decide yourself which points will be (xi,yi)

OpenStudy (anonymous):

umm this confused me more js :S can u explain how u got this? and what it means